OpenStudy (anonymous):

factor the polynomial 2my+7x+7m+2xy

OpenStudy (cwrw238):

= 2my + 7m + 2xy + 7x = m(2y + 7) + x(2y + 7) = (m + x)(2y + 7)

OpenStudy (anonymous):

ty

OpenStudy (cwrw238):

yw